OpenStudy (anonymous):

A football player ran or a gain of 7 1/3 yards. What is 7 1/3 as an improper fraction? 28/3 22/3 20/3 71/3

OpenStudy (zzr0ck3r):

a and b/c = (c*a+b)/c

OpenStudy (anonymous):

simply 22/3

OpenStudy (anonymous):

if you wanted to turn 7 into 3rds \[7=\frac{x}{3}\] how would you solve this for x

OpenStudy (anonymous):

Take the whole number (7) and multiply by 1 in the form of (3/3). \[7 \times \frac{3}{3} = \frac{21}{3}\] Add that to the fraction and you're done. Just keep in mind that every whole number can be represented as a fraction.

OpenStudy (zzr0ck3r):

use the formula I gave you and it will work with all "proper" fractions

OpenStudy (anonymous):

well the question asked for the improper form! and that is 22/3 not 21/3 which is 7

OpenStudy (anonymous):

yes but then you add that to 1/3

OpenStudy (anonymous):

you essentialy have 7 + 1/3 we were simply showing him how to change 7 into 3rds so he can add them together to get 22/3

OpenStudy (anonymous):

you simply gave the answer @chemENGINEER which violates the Code of Conduct which admins take very seriously

OpenStudy (anonymous):

uh oh.

OpenStudy (anonymous):

it's not a big deal, they give you a slap on the wrist. I've given solution manual-esque answers and they've warned me